OpenStudy (anonymous):

Evaluate the expression, using the given value of the variable. z ÷ 2 – 4z + 3(z – 3) when z = 8

OpenStudy (anonymous):

8/2=4 4*8=32 8-3=5 5*3=15 so now you have... 4-32+15 -28+15 -13 I do believe the answer is -13

OpenStudy (anonymous):

i just did the math and the answer is -13 like noneyabusiness said